# Dense Hopfield Networks in the Teacher-Student Setting

### v1

@@ 4.3 Inference temperature vs dataset noise
In the two next Sections, we will discuss the phase diagram when the student is only partially informed about the teacher generative model, i.e. when the Nishimori conditions do not hold. We start with the case where $p=p^{*}$ but $\beta\neq\beta^{*}$, i.e. the inference temperature $T$ is different from the dataset noise $T^{*}$. As we argued in Section 3, the student accurately retrieves $\xi^{*}$ when $T^{*}<T_{\text{crit}}$. On the other hand, we must solve the saddle-points equations (see Eqs. 3) to study $T^{*}>T_{\text{crit}}$.
@@ 4.4 Interaction order and noise tolerance
$\alpha=\frac{Mp!}{N^{p-1}}\ \ \ \hbox{and}\ \ \lambda=\frac{\left[\beta^{*}\right]^{p/2}}{(p/2)!}N^{p/2-1}$
@@
Given $m=0$, the free entropy of the inverse model with $p\geq 3$, $p^{*}=2$ and $\beta=\lambda$ is the same as on the Nishimori line (see Eq. 3 and Appendix D). As a direct consequence, the total entropy is positive outside of the $eR$ phase (see Fig. 6). Additionally, the $p^{*}=2$, $p\geq 3$ phase diagrams with $\beta\neq\lambda$ are identical to the $p=p^{*}$ phase diagrams with $\beta\neq\beta^{*}$, which suggests that $\beta=\lambda$ is optimal for $p^{*}=2$, $p\geq 3$ in the same sense as $\beta=\beta^{*}$ is optimal for $p=p^{*}$ (see Fig. 5). Monte-Carlo simulations confirm that a student with $p\geq 3$ is able to retrieve the pattern of a teacher with $p=2$ and $T^{*}\sim\mathcal{O}\left(N^{1/2}\right)$ (see Fig. 7). However, the $lR$ phase transition is at a higher $T$ in the simulations than on the $\beta=\lambda$ RS phase diagram (see Fig. 5), which means that RSB is necessary to describe it accurately. One could check where replica symmetry holds by evaluating the stability of the RS saddle point throughout the phase diagram.
@@
$\alpha=\frac{M(p/2+1)!}{N^{p/2}}$
is finite. In this finite-noise scaling, $p\geq 3$ requires $\mathcal{O}\left(N^{p/2-1}\right)$ more training examples than $p=2$, which is a lot less than the first scaling. For instance, a student with $p=4$ needs $\mathcal{O}\left(N^{2}\right)$ examples to retrieve $\xi^{*}$. As before, the phase transitions are all first order, the overlap $q^{*}$ stays high throughout the $gR$ and $lR$ phase of $p=4$ and $gR$ is a hard phase. The saddle-point equations (see Eqs. 3) are free from the pattern interference term $\sqrt{\alpha r}x$ present in their $p^{*}=p$ counterparts (see Eqs. 3) until $\beta^{*}$ becomes so small that is approaches $\mathcal{O}\left(N^{2/p-1}\right)$. Therefore, contrary to $p^{*}=p=2$, the network is never in the $SG$ phase. Practically, it means that $p\geq 3$ gives more freedom than $p=2$ for tuning $\beta$ and $\alpha$. The only remaining restriction is that choosing $\alpha$ and $T$ too small puts the network into the inaccurate $eR$ phase resulting from the $kz$ term (see Fig. 6). The saddle point equations can be derived without the RS ansatz because they do not involve $q$ and $r$. Consequently, we expect them to yield an exact solution. Like on the Nishimori line, the total entropy of the paramagnetic phase is always positive, which is consistent with the solution being exact.
@@ 4.5 Robustness against adversarial attacks
Inverse models with $p^{*}=2$ and $p\geq 3$ offer an opportunity to study adversarial attacks in a simple setting because their phase diagrams have regions where the signal retrieval phases ($gR$ and $lR$) overlap with the inaccurate $eR$ phase. Recall that, in the $lR$ phase, a noisy student pattern $\xi$ either converges to $\xi^{*}$ or falls in the paramagnetic phase, depending on the amount of noise that $\xi$ contains initially. The quantity of noise needed to prevent pattern retrieval becomes smaller as one approaches the $lR$ to $P$ phase transition and the basin of attraction of $lR$ shrinks. Similarly, in the region of inaccurate $eR$ where signal retrieval is metastable, patterns $\xi$ that are corrupted by replacing some of their entries $\xi_{i}$ by the components $\sigma^{a}_{i}$ of an example $\sigma^{a}$ may converge to $\sigma^{a}$ when enough entries are replaced. The fraction $\varepsilon$ of entries that need to be replaced becomes smaller as the basin of attraction of inaccurate $eR$ expands and overtakes that of signal retrieval. In practice, an adversary can use this strategy to trick the student into converging to a pattern other than $\xi^{*}$. This scenario is similar to an adversarial attack targeting the input of Krotov’s dense Hopfield network model because the student pattern $\xi$ plays a similar role in the inverse model as the test data in Krotov’s dense Hopfield networks (see Fig. 3, Section 4.2 and Appendix A). In that analogy, the examples $\sigma$ are acting like the neural network weights rather than taking the role of the training data.
We will now investigate what values of the perturbation size $\varepsilon$ are a threat by deriving a formula for the largest $\varepsilon$ such that the student converges to the signal at zero temperature. This largest $\varepsilon$ will be denoted $\varepsilon^{*}$, and we expect it to be a good measure of adversarial robustness. The saddle-point equations with $T=0$ indicate that the student converges to one of the signal retrieval phases if and only if $k<\eta\alpha r^{*}$ (see Eqs. 3). Sampling the initial conditions of $\xi_{i}$ from $\left(1-\varepsilon\right)\delta\left(\xi_{i}-\xi^{*}_{i}\right)+\varepsilon\ \delta\left(\xi_{i}-\sigma^{a}_{i}\right)$ with $\varepsilon\in\left[0,1\right]$, we get
@@
$r^{*}$ $\approx p\left(1-\varepsilon\right)^{p-1}$
@@
$\varepsilon^{p-1}$ $<\eta\alpha\left(1-\varepsilon\right)^{p-1}$
@@
In other terms, the inverse model with $p^{*}=2$ and even $p\geq 3$ is resistant to adversarial attacks of size $\varepsilon^{*}=\frac{\left[\eta\alpha\right]^{\frac{1}{p-1}}}{\left[\eta\alpha\right]^{\frac{1}{p-1}}+1}$ and smaller. For $p=4$, $\varepsilon^{*}$ is in good agreement with Monte-Carlo simulations of the inverse model corrupted by a typical example (see Fig. 8). This comparison is good evidence that our solution of the finite-noise scaling is indeed exact. Additionally, $\varepsilon^{*}$ is a decent approximation of empirical robustness even when the inverse model is corrupted by the example that has the largest overlap with $\xi^{*}$. Just like adversarial attacks targeting more complicated neural networks [40, 41], our example-based attack can be hard to detect at low $\varepsilon$ because a few adversarially perturbed entries $\xi_{i}$ do not look very different from a low amount of meaningless noise.
@@
Figure 8: Monte-Carlo simulations of the overlap $q^{*}$ as a function of $\alpha$ and $\varepsilon$ in the inverse model with $p^{*}=2$, $\beta^{*}=1-\frac{1}{\sqrt{2}}$, $p=4$, $\beta=\infty$ and $N=1024$. The simulation results are averaged over $L=100$ student patterns. On the left plot, the inverse model is corrupted by an example $\sigma^{a}$ that has a small overlap with $\xi^{*}$ in absolute value. On the right plot, it is corrupted by the example that has the largest overlap with $\xi^{*}$. The black line $\varepsilon^{*}=\frac{\alpha^{1/3}}{\alpha^{1/3}+1}$ is our analytical formula for the largest adversarial perturbation $\varepsilon$ such that the student retrieves $\xi^{*}$ rather than the example $\sigma^{a}$.
@@
Interestingly, $\varepsilon^{*}$ grows monotonically with $\alpha$, which is consistent with the common observation that larger neural networks are also more adversarially robust [43, 60, 61, 62, 63, 64, 65]. At first glance, this effect can be counter-intuitive because adversarial vulnerability looks like a form of overfitting [42]. In our model, however, all examples work together to stabilize the $lR$ phase, and the best way to push the student into the $eR$ phase is to perturb it with a single example. Therefore, it is not surprising that increasing $\alpha$ makes the student more robust. We recall that the examples $\sigma$ are a feature-based representation of $\xi^{*}$. Interestingly, it means that the underlying mechanism of our example-based attack is conceptually similar to gradient-based attacks targeting many common types of neural networks [42]. In fact, gradient-based attacks find features stored in neural network weights and add them to the data in order to fool the network [42, 66, 67, 68]. It would be interesting to investigate, both empirically and theoretically, if only a small number of weights are involved in constructing these adversarial attacks. If it is the case, it could explain why larger neural networks are often more robust. In general, we expect this kind of one-example attack to be possible in any region of signal retrieval that overlaps with the inaccurate $eR$ phase. Using $p\neq p^{*}$ may not be a necessary ingredient of adversarial vulnerability in more general models with other sources of mismatch, but in our case it ensures that the signal retrieval phases intersect the inaccurate $eR$ phase. Conversely, the accurate $eR$ phase is by definition robust to adversarial attacks since retrieving an example $\sigma^{a}$ is the same as recovering $\xi^{*}$. This distinction clarifies why the dense Hopfield networks designed by Krotov are adversarially robust in the prototype phase despite being adversarially vulnerable in the feature phase. In fact, Krotov observed that adversarial attacks are unsuccessful in the prototype phase specifically because they retrieve stored examples that are semantically meaningful [37]. In summary, our model yields two main results concerning adversarial examples. First of all, it suggests a reason why large feature-based neural networks are more adversarially robust than smaller ones. Second of all, it clarifies why dense Hopfield networks are much more robust in the prototype phase than in the feature phase.
@@ 5 Conclusion
Dense networks with exponential interactions have been argued to be the $p\rightarrow\infty$ limit of the $p$-body models [69]. It would be interesting to see if they can achieve $\mathcal{O}\left(N\right)$ noise tolerance at the cost of an exponential number of training examples. More generally, studying exponential models in the teacher-student setting would be an interesting extension of this work and could be used to complement existing studies of the direct model [69, 70]. A caveat of our model is that the teacher has only one pattern. In fact, we would need to use a teacher with at least two patterns to describe more completely the kind of adversarial attack aiming to misclassify data. It should be possible to study this kind of teacher by using an approach similar to [71]. On the practical side, we highlight the untapped benefits of using $p$-body models to either resist an extensive amount of noise in the feature phase or improve adversarial robustness in the prototype phase. Overall, we stress that further investigations of dense Hopfield networks could unlock their true potential.
@@ Acknowledgements
This work was partially supported by project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU. The work was also supported by the project PRIN22TANTARI ”Statistical Mechanics of Learning Machines: from algorithmic and information-theoretical limits to new biologically inspired paradigms” 20229T9EAT – CUP J53D23003640001. DT also acknowledges GNFM-Indam.
Appendix A Gardner’s Hamiltonian vs Krotov’s Hamiltonian
Consider the generalized Hopfield Hamiltonian $H\left[\sigma|\xi\right]=-\sum_{i_{1}<...<i_{p}=1}^{N}J_{i_{1}...i_{p}}\sigma_{i_{1}}...\sigma_{i_{p}}$ with $p$-body interactions $J_{i_{1}...i_{p}}=\frac{p!}{N^{p-1}}\sum_{\mu=1}^{M}\xi_{i_{1}}^{\mu}...\xi_{i_{p}}^{\mu}$ described by Gardner [30], where $M$ indicates the number of patterns $\xi^{\mu}$ used to construct $J$, and $N$ denotes the number of components of each pattern $\xi^{\mu}$ and example $\sigma$. In this Section, we will omit $\xi$ in the argument of $H\left[\sigma|\xi\right]$ and write $H\left[\sigma\right]$ instead for notational simplicity. Unless indicated otherwise, we will assume a large number number of components $N\gg 1$ and patterns $M\sim\mathcal{O}\left(N^{p-1}\right)$. We will start by comparing it to the dense Hopfield network Hamiltonian $\mathcal{H}\left[\sigma\right]=-\frac{1}{N^{p-1}}\sum_{\mu}\left(\sum_{i}\xi_{i}^{\mu}\sigma_{i}\right)^{p}$ studied by Krotov [26].

### v2

@@ 4.4 Inference temperature vs dataset noise
In the two next Sections, we will discuss the phase diagram when the student is only partially informed about the teacher generative model, i.e. when the Nishimori conditions do not hold. We start with the case where $p=p^{*}$ but $\beta\neq\beta^{*}$, i.e. the inference temperature $T$ is different from the dataset noise $T^{*}$. As we argued in Section 3.1, the student accurately retrieves $\xi^{*}$ when $T^{*}<T_{\text{crit}}$. On the other hand, we must solve the saddle-points equations (see Eqs. 3.1) to study $T^{*}>T_{\text{crit}}$.
@@ 4.5.1 Large noise scaling
$\alpha=\frac{Mp!}{N^{p-1}}\,,\quad\hbox{and}\quad\lambda=\frac{\left[\beta^{*}\right]^{p/2}}{(p/2)!}N^{p/2-1}\,,$
@@
Given $m=0$, the free entropy of the inverse model with $p\geq 3$, $p^{*}=2$ and $\beta=\lambda$ is the same as on the Nishimori line (see Eq. 3.1 and Appendix D). As a direct consequence, the total entropy is positive outside of the $eR$ phase (see Fig. 6). Additionally, the $p^{*}=2$, $p\geq 3$ phase diagrams with $\beta\neq\lambda$ are identical to the $p=p^{*}$ phase diagrams with $\beta\neq\beta^{*}$, which suggests that $\beta=\lambda$ is optimal for $p^{*}=2$, $p\geq 3$ in the same sense as $\beta=\beta^{*}$ is optimal for $p=p^{*}$ (see Fig. 5). Monte-Carlo simulations confirm that a student with $p\geq 3$ is able to retrieve the pattern of a teacher with $p=2$ and $T^{*}\sim\mathcal{O}\left(N^{1/2}\right)$ (see Fig. 7). However, the $lR$ phase transition is at a higher $T$ in the simulations than on the $\beta=\lambda$ RS phase diagram (see Fig. 5), which means that RSB is necessary to describe it accurately. One could check where replica symmetry holds by evaluating the stability of the RS saddle point throughout the phase diagram.
4.5.2 Finite noise scaling
@@
$\alpha=\frac{M(p/2+1)!}{N^{p/2}}\,,$
is finite. In this finite-noise scaling, $p\geq 3$ requires $\mathcal{O}\left(N^{p/2-1}\right)$ more training examples than $p=2$, which is a lot less than the first scaling. For instance, a student with $p=4$ needs $\mathcal{O}\left(N^{2}\right)$ examples to retrieve $\xi^{*}$. As before, the phase transitions are all first order, the overlap $q^{*}$ stays high throughout the $gR$ and $lR$ phase of $p=4$ and $gR$ is a hard phase. The saddle-point equations (see Eqs. 3.2) are free from the pattern interference term $\sqrt{\alpha r}x$ present in their $p^{*}=p$ counterparts (see Eqs. 3.1) until $\beta^{*}$ becomes so small that is approaches $\mathcal{O}\left(N^{2/p-1}\right)$. Therefore, contrary to $p^{*}=p=2$, the network is never in the $SG$ phase. Practically, it means that $p\geq 3$ gives more freedom than $p=2$ for tuning $\beta$ and $\alpha$. The only remaining restriction is that choosing $\alpha$ and $T$ too small puts the network into the inaccurate $eR$ phase resulting from the $kz$ term (see Fig. 6). The saddle point equations can be derived without the RS ansatz because they do not involve $q$ and $r$. Consequently, we expect them to yield an exact solution. Like on the Nishimori line, the total entropy of the paramagnetic phase is always positive, which is consistent with the solution being exact.
@@ 4.6 Robustness against adversarial attacks
Inverse models with $p^{*}=2$ and $p\geq 3$ offer an opportunity to study adversarial attacks in a simple setting because their phase diagrams have regions where the signal retrieval phases ($gR$ and $lR$) overlap with the inaccurate $eR$ phase. Recall that, in the $lR$ phase, a noisy student pattern $\xi$ either converges to $\xi^{*}$ or falls in the paramagnetic phase, depending on the amount of noise that $\xi$ contains initially. The quantity of noise needed to prevent pattern retrieval becomes smaller as one approaches the $lR$ to $P$ phase transition and the basin of attraction of $lR$ shrinks. Similarly, in the region of inaccurate $eR$ where signal retrieval is metastable, patterns $\xi$ that are corrupted by replacing some of their entries $\xi_{i}$ by the components $\sigma^{a}_{i}$ of an example $\sigma^{a}$ may converge to $\sigma^{a}$ when enough entries are replaced. The fraction $\varepsilon$ of entries that need to be replaced becomes smaller as the basin of attraction of inaccurate $eR$ expands and overtakes that of signal retrieval. In practice, an adversary can use this strategy to trick the student into converging to a pattern other than $\xi^{*}$. This scenario is similar to an adversarial attack targeting the input of K & H’s dense Hopfield network model because the student pattern $\xi$ plays a similar role in the inverse model as the test data in K & H’s dense Hopfield networks (see Fig. 3, Section 4.3 and Appendix A). In that analogy, the examples $\sigma$ are acting like the neural network weights rather than taking the role of the training data.
We will now investigate what values of the perturbation size $\varepsilon$ are a threat by deriving a formula for the largest $\varepsilon$ such that the student converges to the signal at zero temperature. This largest $\varepsilon$ will be denoted $\varepsilon^{*}$, and we expect it to be a good measure of adversarial robustness. The saddle-point equations with $T=0$ indicate that the student converges to one of the signal retrieval phases if and only if $k<\eta\alpha r^{*}$ (see Eqs. 3.2). Sampling the initial conditions of $\xi_{i}$ from $\left(1-\varepsilon\right)\delta\left(\xi_{i}-\xi^{*}_{i}\right)+\varepsilon\ \delta\left(\xi_{i}-\sigma^{a}_{i}\right)$ with $\varepsilon\in\left[0,1\right]$, we get
@@
$r^{*}$ $\approx p\left(1-\varepsilon\right)^{p-1}\,,$
@@
$\varepsilon^{p-1}$ $<\eta\alpha\left(1-\varepsilon\right)^{p-1}\,,$
@@
In other terms, the inverse model with $p^{*}=2$ and even $p\geq 3$ is resistant to adversarial attacks of size $\varepsilon^{*}=\frac{\left[\eta\alpha\right]^{\frac{1}{p-1}}}{\left[\eta\alpha\right]^{\frac{1}{p-1}}+1}$ and smaller. For $p=4$, $\varepsilon^{*}$ is in good agreement with Monte-Carlo simulations of the inverse model corrupted by a typical example (see Fig. 8). This comparison is good evidence that our solution of the finite-noise scaling is indeed exact. Additionally, $\varepsilon^{*}$ is a decent approximation of empirical robustness even when the inverse model is corrupted by the example that has the largest overlap with $\xi^{*}$. A similar construction with the perturbation sampled uniformly at random gives $k\sim\mathcal{O}\left(N^{1/2-p/2}\right)\approx 0$, so adversarial attacks are much more efficient at fooling the model than random noise. Just like adversarial attacks targeting more complicated neural networks [40, 41], our example-based attack can be hard to detect at low $\varepsilon$ because a few adversarially perturbed entries $\xi_{i}$ do not look very different from a low amount of meaningless noise.
@@
Figure 8: Monte-Carlo simulations of the overlap $q^{*}$ as a function of $\alpha$ and adversarial attack size $\varepsilon$ in the inverse model with $p^{*}=2$, $\beta^{*}=1-\frac{1}{\sqrt{2}}$, $p=4$, $\beta=\infty$ and $N=1024$. The simulation results are averaged over $L=100$ student patterns. On the left plot, the inverse model is corrupted by an example $\sigma^{a}$ that has a small overlap with $\xi^{*}$ in absolute value. On the right plot, it is corrupted by the example that has the largest overlap with $\xi^{*}$. The black line $\varepsilon^{*}=\frac{\alpha^{1/3}}{\alpha^{1/3}+1}$ is our analytical formula for the largest adversarial perturbation $\varepsilon$ such that the student retrieves $\xi^{*}$ rather than the example $\sigma^{a}$.
@@
Moreover, $\varepsilon^{*}$ grows monotonically with $\alpha$, which is consistent with the common observation that larger neural networks are also more adversarially robust [43, 66, 67, 68, 69, 70, 71]. At first glance, this effect can be counter-intuitive because adversarial vulnerability looks like a form of overfitting [42]. In our model, however, all examples work together to stabilize the $lR$ phase, and the best way to push the student into the $eR$ phase is to perturb it with a single example. Therefore, it is not surprising that increasing $\alpha$ makes the student more robust. We recall that the examples $\sigma$ are a feature-based representation of $\xi^{*}$. Interestingly, it means that the underlying mechanism of our example-based attack is conceptually similar to gradient-based attacks targeting many common types of neural networks [42]. In fact, gradient-based attacks find features stored in neural network weights and add them to the data in order to fool the network [42, 72, 73, 74]. It would be interesting to investigate, both empirically and theoretically, if only a small number of weights are involved in constructing these adversarial attacks. If it is the case, it could explain why larger neural networks are often more robust. In general, we expect this kind of one-example attack to be possible in any region of signal retrieval that overlaps with the inaccurate $eR$ phase. Using $p\neq p^{*}$ may not be a necessary ingredient of adversarial vulnerability in more general models with other sources of mismatch, but in our case it ensures that the signal retrieval phases intersect the inaccurate $eR$ phase. Conversely, the accurate $eR$ phase is by definition robust to adversarial attacks since retrieving an example $\sigma^{a}$ is the same as recovering $\xi^{*}$. This distinction clarifies why the dense Hopfield networks designed by K & H are adversarially robust in the prototype phase despite being adversarially vulnerable in the feature phase. In fact, K & H observed that adversarial attacks are unsuccessful in the prototype phase specifically because they retrieve stored examples that are semantically meaningful [37]. In summary, our model yields two main results concerning adversarial examples. First of all, it suggests a reason why large feature-based neural networks are more adversarially robust than smaller ones. Second of all, it clarifies why dense Hopfield networks are much more robust in the prototype phase than in the feature phase.
@@ 5 Conclusion
Dense networks with exponential interactions have been argued to be the $p\rightarrow\infty$ limit of the $p$-body models [75]. It would be interesting to see if they can achieve $\mathcal{O}\left(N\right)$ noise tolerance at the cost of an exponential number of training examples. More generally, studying exponential models in the teacher-student setting would be an interesting extension of this work and could be used to complement existing studies of the direct model [75, 76]. A caveat of our model is that the teacher has only one pattern. In fact, we would need to use a teacher with at least two patterns to describe more completely the kind of adversarial attack aiming to misclassify data. It should be possible to study this kind of teacher by using an approach similar to [77]. In particular, [16] and [77] argue that the performance of restricted Boltzmann machines with a finite number $P$ of i.i.d. planted patterns is independent of $P$ in the teacher-student setting. It would be interesting to investigate whether this characteristic also holds for $p$-body dense networks. On the practical side, we highlight the untapped benefits of using $p$-body models to either resist an extensive amount of noise in the feature phase or improve adversarial robustness in the prototype phase. Overall, we stress that further investigations of dense Hopfield networks could unlock their true potential.
@@ Acknowledgements
Funding information
This work was partially supported by project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU. The work was also supported by the project PRIN22TANTARI "Statistical Mechanics of Learning Machines: from algorithmic and information-theoretical limits to new biologically inspired paradigms" 20229T9EAT – CUP J53D23003640001. DT also acknowledges GNFM-Indam.
Code availability
The figures can be reproduced using the code available on this public Github repository.
Appendix A Gardner’s Hamiltonian vs K & H’s Hamiltonian
Consider the generalized Hopfield Hamiltonian $H\left[\sigma|\xi\right]=-\sum_{i_{1}<...<i_{p}=1}^{N}J_{i_{1}...i_{p}}\sigma_{i_{1}}...\sigma_{i_{p}}$ with $p$-body interactions $J_{i_{1}...i_{p}}=\frac{p!}{N^{p-1}}\sum_{\mu=1}^{M}\xi_{i_{1}}^{\mu}...\xi_{i_{p}}^{\mu}$ described by Gardner [30], where $M$ indicates the number of patterns $\xi^{\mu}$ used to construct $J$, and $N$ denotes the number of components of each pattern $\xi^{\mu}$ and example $\sigma$. In this Section, we will omit $\xi$ in the argument of $H\left[\sigma|\xi\right]$ and write $H\left[\sigma\right]$ instead for notational simplicity. Unless indicated otherwise, we will assume a large number number of components $N\gg 1$ and patterns $M\sim\mathcal{O}\left(N^{p-1}\right)$. We will start by comparing it to the dense Hopfield network Hamiltonian $\mathcal{H}\left[\sigma\right]=-\frac{1}{N^{p-1}}\sum_{\mu}\left(\sum_{i}\xi_{i}^{\mu}\sigma_{i}\right)^{p}$ studied by K & H [26].